OpenStudy (rayford):

solve sqrt 2x+15=sqrt 5

OpenStudy (anonymous):

\[\sqrt{2x+15}=\sqrt{5}\] is this the question???

OpenStudy (rayford):

yes :)

OpenStudy (anonymous):

if it is......then by squaring both the sides we arrive at.... \[(\sqrt{2x+15})^{2}=(\sqrt{5})^{2}\] \[2x+15\] \[2x=-10\] \[x=-5\]

OpenStudy (anonymous):

the second equation is.... \[2x+15=5\]

OpenStudy (rayford):

thank you very much then you plug it back in and it equals one :).